Find the matrix A for T relative to the basis B. T: R2 + R2, T(x, y) = (3x - y, 4x), B = {(-2, 1), (-1, 1)} A =Let B = {(1, 3), (-2,-2)} and B = {(-12, 0), (-4,4)} be bases for R2, and let 0 2 A = 3 4 be the matrix for T: R2 + R2 relat(d) Find [T(v)]B two ways. [T(v)]g = p-1[T(v)]B [T(v)]B = A[v]B
Answer #1

Given:
T: R2 R and  

T(2, y) = (5.r - y, 4.1)

B = {(-2,1),(-1,1)}

Let, (5.1 – y, 4.0) = c(-2,1) + c(-1,1)

→ 5.r - y = (-2) + (-1)02: 4.c = c + c

5.6 - +4.x = (-2)c1 + (-1) c2 + 1 + Cm

9.0 - 4 -C1

→ C1 -9.7 +y

C2 = 4.0 - (-9.2 + y)

C2 13.C - Y

..C1 -9.r + y; C2 13.0 Y

Hence, T(2, y) = (-9.r + y) (-2,1) + (13.1 - y) (-1,1)

\Rightarrow T(-2,1)=(-9\times (-2)+1)(-2,1)+ (13\times (-2)-1)(-1,1)

      ( 18 -11-21)+(-26-11-1.1)

197-2.11-27(-1.1)

and T(-1,1)= (-9% (-1) + 1) (-2,1) + (13 x (-1) - 1)(-1,1)

=(9+1)(-2,1)+ (-13-1)(-1,1)

={\color{Blue} 10}(-2,1){\color{Blue} -14}(-1,1)

:. A = [T]B = 19 10 -27 -14

which is the required matrix A' for T relative to the basis B'

Due to HOMEWORKLIB POLICY we are required to do just the first question please ask the second question separately.

Hope this helps!